# Data-Efficient Kernel Methods for Learning Differential Equations and Their Solution Operators: Algorithms and Error Analysis

### v1

@@ Abstract.
We introduce a novel kernel-based framework for learning differential equations and their solution maps that is efficient in data requirements, in terms of solution examples and amount of measurements from each example, and computational cost, in terms of training procedures. Our approach is mathematically interpretable and backed by rigorous theoretical guarantees in the form of quantitative worst-case error bounds for the learned equation. Numerical benchmarks demonstrate significant improvements in computational complexity and robustness while achieving one to two orders of magnitude improvements in terms of accuracy compared to state-of-the-art algorithms.
@@
We present a novel algorithm inspired by kernel methods and Gaussian processes for learning differential equations and their solution operators in scarce data regimes. Our approach: (a) is significantly more efficient than state-of-the-art methods, including neural networks, in terms of required data and computational time. In fact, we obtain one to two orders of magnitude improvement in accuracy on a number of benchmarks; (b) is supported by rigorous theory featuring the first quantitative worst-case error bounds for equation learning; and (c) can solve previously intractable scientific computing problems such as one-shot operator learning and learning of variable-coefficient PDEs in extremely scarce data regimes.
@@ 2. The proposed method
with a smooth coefficient $a:[0,1]\mapsto(0,+\infty]$. Writing $y=(t,x)$ and letting $\mathcal{Y}=(0,T]\times(0,1)$ and introducing the differential operators $L_{1}:u\mapsto u$, $L_{2}:u\mapsto\partial_{t}u$, $L_{3}:u\mapsto\partial_{x}u$, and $L_{4}:u\mapsto\partial_{xx}u$ we can cast (3) in the form (2) by writing $\mathscr{P}(u)(t,x)=P(t,x,u(t,x),L_{2}u(t,x),L_{3}u(t,x),L_{4}u(t,x))$ with the nonlinear map
@@ 3.2. Quantitative error bounds
We emphasize that the above error bounds are quite strong, and give pointwise control over the errors of the $\widehat{u}^{m}_{M,N}$ and $\widehat{P}_{M,N}$ over subsets of their domains provided that the training data is sufficiently space-filling. This is in line with the theory of scattered data approximation and suggests various avenues for experimental design of equation learning algorithms. Furthermore, the bound for $\widehat{P}_{M,N}$ implies that the learned PDE is accurate for new input-output pairs $(u,f)$ that are close to the training data in an appropriate sense, i.e., if $\Phi(u,\cdot)\in B$. In Section B.3 we use this observation to extend the error bound for $\widehat{P}_{M,N}$ to the corresponding pesudo inverse $\widehat{\mathscr{P}}_{M,N}^{\dagger}$.
@@ 4.2. Implementing 1-step KEqL and its reduced version
$\widehat{u}^{m}(y)=\mathsf{U}(\phi,y)^{T}\widehat{\alpha}^{m},\quad\widehat{P}(s)=\mathsf{P}(\widehat{S},s)^{T}\widehat{\beta},$
@@
(14) $(\widehat{\boldsymbol{\alpha}},\widehat{\beta})=$ $\operatorname*{arg\,min}_{\boldsymbol{\alpha}\in(\mathbb{R}^{QK})^{M},\>\beta\in\mathbb{R}^{MK}}$ $\beta^{T}\mathsf{P}(S(\boldsymbol{\alpha}),S(\boldsymbol{\alpha})\beta+\lambda\sum_{m=1}^{M}(\alpha^{m})^{T}\mathsf{U}(\phi,\phi)\alpha^{m},$
@@
where $S(\boldsymbol{\alpha})=\cup_{m=1}^{M}S^{m}(\alpha^{m}),\quad\text{and}\quad S^{m}(\alpha^{m})=\Phi(U(\phi,\cdot)^{T}\alpha^{m},Y).$
Here we introduced the block vector $\boldsymbol{\alpha}:=(\alpha^{1},\dots,\alpha^{M})\in(\mathbb{R}^{QK})^{M}$ for coefficient vectors $\alpha^{m}\in\mathbb{R}^{QK}$, along with the matrices $\mathsf{U}(\phi,Y^{m})\in\mathbb{R}^{QK\times N}$ with columns $\mathsf{U}(\phi,y^{m}_{n})$, and $\mathsf{P}(S,S^{m})\in\mathbb{R}^{MK\times K}$ with columns $\mathsf{P}(S,s^{m}_{k})$.
@@ 5.3.3. Results
We begin by considering the filtering and equation learning errors for one-shot learning of Burgers’. These results are show in the top row of Figure 3(A) where we are reporting the errors for a fixed smooth initial condition with different number of observation points. Note, these errors were computed over the training data but they are different from the training error of the algorithm as they compare the filtered solution and the learned PDE on the test mesh $Y_{\text{test}}$. Since the observation points are random we are reporting the average errors along with the best and worst errors over different runs giving the shaded regions. We found that 1-step KEqL leads to the best filtering errors with a wider performance gap when $N_{\mathcal{Y}}$ is small. Interestingly, the performance gap appears to be more pronounced in the equation learning case where we observed an order of magnitude improvement over SINDy and almost two orders improvement over PINN-SR. We observed significantly larger errors for PINN-SR here which we attribute to difficulties in tuning and optimizing PINN models. In fact, we could often bring these errors down by hand tuning the algorithm in each instance of the experiments but automatic strategies proposed in [21] appeared to be very sensitive to the location of observation points. This issue was exacerbated further when experimenting with randomized initial conditions; see SI D.
In Figure 3(B, C) we show two enlightening examples that show the difference in performance of 1-step methods vs 2-step. In (B) we hand picked an initial condition that leads to a solution with multiple (near) shocks. We see that despite scarcity of observations, 1-step KEqL and PINN-SR are able to capture the general shape of the solution while 2-step methods are significantly worse. In (C) we show an instance where we only observed the value of the solution on the boundary of the domain while in row (D) we solved the learned equation for a new initial condition. Once again we see that 1-step methods give better performance for filtering, however, when solving the learned PDE a noticable performance gap appears across the three methods with 1-step KEqL giving the best solution. This example further reflects the performance gaps observed in panel (A), with the equation learning performance gaps being more pronounced compared to filtering.
@@
Figure 3. Representative numerical results for Burgers’ PDE (19): (A) The filtering and equation learning errors computed for the training functions for 1-step KEqL, SINDy, and PINN-SR using $M=1$ training pairs with different number of interior observations $N_{\mathcal{Y}}$; (B) An example application for an initial condition that leads to multiple shocks with scarce observations depicting the quality of filtering obtained using 1-step and 2-step methods; (C) Similar setup to row (B) with a smooth solution that is only observed on the boundary; (D) Depicting the solution to the PDEs that were learned in row (C) for a new initial condition.
@@ 6. Discussion
In conclusion, we presented the 1- and 2-step formulations of KEqL as an algorithm for learning nonlinear PDEs as well as their solution operators and filtering of observed states in scarce data regimes. Our theoretical results provided quantitative error bounds and convergence rates for our algorithms while our numerical experiments demonstrated significant gains in accuracy compared to existing methods in the literature. Additionally, our exposition unifies various problems of interest to scientific machine learning under the same umbrella, i.e., equation learning, operator learning, and PDE solvers, all viewed as optimal recovery problems.
@@
JFOR, AH and BH were supported by the National Science Foundations grants DMS-2208535 ”Machine Learning for Inverse Problems” and DMS-2337678 ”Gaussian Processes for Scientific Machine Learning: Theoretical Analysis and Computational Algorithms”. YJ and HO acknowledge support from the Air Force Office of Scientific Research under MURI awards number FA9550-20-1-0358 (Machine Learning and Physics-Based Modeling and Simulation), FOA-AFRL-AFOSR-2023-0004 (Mathematics of Digital Twins), by the Department of Energy under award number DE-SC0023163 (SEA-CROGS: Scalable, Efficient, and Accelerated Causal Reasoning Operators, Graphs and Spikes for Earth and Embedded Systems) and by the National Science Foundations under award number 2425909 (Discovering the Law of Stress Transfer and Earthquake Dynamics in a Fault Network using a Computational Graph Discovery Approach). Additionally, HO acknowledges support from the DoD Vannevar Bush Faculty Fellowship Program.
@@ B.4.2. Representer theorem for 1-step KEqL (Proof of Theorem 2)
To simplify notation we first prove the theorem with $M=1$, i.e., with training data $(u(Y^{1}),f)$. Let us recall the corresponding optimal recovery problem for convenience:
@@
where $S=\{s_{1},\dots,s_{K}\}$, and $s_{k}:=\Phi(v,y_{k})$ for $y_{k}\in Y$ where we recall that $Y$ is the dense set of collocation points that we use to impose the PDE constraint.
@@
$\widehat{u}(y)=\mathsf{U}(\phi,y)^{T}\widehat{\alpha},\quad\widehat{P}(s)=\mathsf{P}(\widehat{S},s)^{T}\widehat{\beta},$
@@
(32) $(\widehat{\alpha},\widehat{\beta})=$ $\operatorname*{arg\,min}_{\alpha\in\mathbb{R}^{QK},\>\beta\in\mathbb{R}^{K}}$ $\beta^{T}\mathsf{P}(S,S)\beta+\lambda_{1}\alpha^{T}\mathsf{U}(\phi,\phi)\alpha,$
s.t. $\mathsf{U}(\phi,Y^{1})^{T}\alpha=u(Y^{1}),\quad\text{and}\quad\mathsf{P}(S,S)^{T}\beta=f(Y).$
where $S=\{s_{1},\dots s_{K}\}$ but now each $s_{k}=\left(y_{k},\mathsf{U}(\phi,\phi^{1}_{k})^{T}\alpha,\dots,\mathsf{U}(\phi,\phi^{Q}_{k})^{T}\alpha\right)$. Recall our notation $\mathsf{U}(\phi,Y^{m})\in\mathbb{R}^{QK\times N}$ for the matrix with columns $\mathsf{U}(\phi,y_{n})$, and $\mathsf{P}(S,S)\in\mathbb{R}^{K\times K}$ for the matrix with columns with columns $\mathsf{P}(S,s_{k})$.
@@
For the general proof for an arbitrary number of functions $M$ the same idea follows since the equivalent form Equation 33 now contains the slack tensor $Z\in\mathbb{R}^{K\times Q\times M}$ instead of the matrix $Z$ above since each $v^{m}$ will be represented by a matrix $Z^{m}$ in our proof. Additionally $S$ will then contain not only $K$ points but $KM$ points corresponding to each of the $v^{m}$ and their requisite partial derivatives evaluated at the collocation points.
@@ D.2.1. Experiments with increasing number of observations
We analyzed the filtering and equation learning errors for one-shot learning of Burgers’ equation with different numbers of observation points and different ICs. The results are presented in Figure 8, where we report the errors for different smooth initial conditions. Since both the observation points and initial conditions are chosen randomly, we report the average errors along with the best and worst errors across multiple runs. Convergence history of the optimization algorithm for some sample cases of learning the equation is shown in Figure 11(A-B).
A similar conclusion can be drawn here as in Figure 3(A), where the 1-step KEqL method continues to achieve the lowest filtering and equation learning errors, with a larger performance gap when $N_{\mathcal{Y}}$ is small. However, we now observe that the KEqL and SINDy methods exhibit greater robustness across different experiments with varying initial conditions, whereas the performance of the PINN-SR method deteriorates. This decline is reflected in the increased variance in the results, which we attribute to the use of the same model across different experiments. This suggests that PINN-SR may require tuning for each specific set of observations and initial conditions, whereas such tuning was not necessary for the other two methods.
[…]
Figure 8. The filtering and equation learning errors computed for the training functions for 1-step KEqL, SINDy, and PINN-SR using $M=1$ training pairs with different number of interior observations $N_{\mathcal{Y}}$ for Burgers’ PDE (19) with different ICs.
@@ D.2.3. A one-shot learning experiment
Figure 10. Quantitative values of relative filtering and operator learning errors for Burgers’ PDE Equation 19 with different ICs corresponding to the one-shot learning examples Ex. 1 (Section D.2.2) and Ex. 2 (Section D.2.3) in Figure 3(B-D).

### v2

@@ Abstract.
We introduce a novel kernel-based framework for learning differential equations and their solution maps that is efficient in data requirements, in terms of solution examples and amount of measurements from each example, and computational cost, in terms of training procedures. Our approach is mathematically interpretable and backed by rigorous theoretical guarantees in the form of quantitative worst-case error bounds for the learned equation. Numerical benchmarks demonstrate significant improvements in computational complexity and robustness while achieving one to two orders of magnitude improvements in terms of accuracy compared to state-of-the-art algorithms. In comparison to equivalent neural net methods, our approach is significantly more robust to the choice of hyperparameters and does not require close human supervision during training.
@@
We present a novel algorithm inspired by kernel methods and Gaussian processes for learning differential equations and their solution operators in scarce data regimes. Our approach: (a) is significantly more efficient than state-of-the-art methods, including neural networks, in terms of required data and computational time. In fact, we obtain one to two orders of magnitude improvement in accuracy on a number of benchmarks; (b) is significantly more robust to choice of hyperparameters and does not require close human supervision during training in comparison to equivalent neural net models; (c) is supported by rigorous theory featuring the first quantitative worst-case error bounds for equation learning; and (d) can solve previously intractable scientific computing problems such as one-shot operator learning and learning of variable-coefficient PDEs in extremely scarce data regimes.
@@ 2. The proposed method
with a smooth coefficient $a:[0,1]\to(0,+\infty]$. Writing $y=(t,x)$ and letting $\mathcal{Y}=(0,T]\times(0,1)$ and introducing the differential operators $L_{1}:u\mapsto u$, $L_{2}:u\mapsto\partial_{t}u$, $L_{3}:u\mapsto\partial_{x}u$, and $L_{4}:u\mapsto\partial_{xx}u$ we can cast (3) in the form (2) by writing $\mathscr{P}(u)(y)=P(t,x,u(t,x),L_{2}u(t,x),L_{3}u(t,x),L_{4}u(t,x))$ with the nonlinear map
@@ 3.2. Quantitative error bounds
We emphasize that the above error bounds are quite strong, and give pointwise control over the errors of the $\widehat{u}^{m}_{M,N}$ and $\widehat{P}_{M,N}$ over subsets of their domains provided that the training data is sufficiently space-filling. This is in line with the theory of scattered data approximation and suggests various avenues for experimental design of equation learning algorithms. Furthermore, the bound for $\widehat{P}_{M,N}$ implies that the learned PDE is accurate for new input-output pairs $(u,f)$ that are close to the training data in an appropriate sense, i.e., if $\Phi(u,\cdot)\in B$. In Section B.3 we use this observation to extend the error bound for $\widehat{P}_{M,N}$ to the corresponding pseudo inverse $\widehat{\mathscr{P}}_{M,N}^{\dagger}$.
@@ 4.2. Implementing 1-step KEqL and its reduced version
$\widehat{u}^{m}(y)=\mathsf{U}(\phi,y)^{T}\widehat{\alpha}^{m},\quad\widehat{P}(s)=\mathsf{P}(S(\widehat{\boldsymbol{\alpha}}),s)^{T}\widehat{\beta},$
@@
(14) $(\widehat{\boldsymbol{\alpha}},\widehat{\beta})=$ $\operatorname*{arg\,min}_{\boldsymbol{\alpha}\in(\mathbb{R}^{QK})^{M},\>\beta\in\mathbb{R}^{MK}}$ $\beta^{T}\mathsf{P}(S(\boldsymbol{\alpha}),S(\boldsymbol{\alpha}))\beta+\lambda\sum_{m=1}^{M}(\alpha^{m})^{T}\mathsf{U}(\phi,\phi)\alpha^{m},$
@@
where $S(\boldsymbol{\alpha})=\cup_{m=1}^{M}S^{m}(\alpha^{m}),\quad\text{and}\quad S^{m}(\alpha^{m})=\Phi(\mathsf{U}(\phi,\cdot)^{T}\alpha^{m},Y).$
Here we introduced the block vector $\boldsymbol{\alpha}:=(\alpha^{1},\dots,\alpha^{M})\in(\mathbb{R}^{QK})^{M}$ for coefficient vectors $\alpha^{m}\in\mathbb{R}^{QK}$, along with the matrices $\mathsf{U}(\phi,Y^{m})\in\mathbb{R}^{QK\times N}$ with columns $\mathsf{U}(\phi,y^{m}_{n})$, and $\mathsf{P}(S(\boldsymbol{\alpha}),S^{m}(\alpha^{m}))\in\mathbb{R}^{MK\times K}$ with columns $\mathsf{P}(S,s^{m}_{k})$.
@@ 5.3.3. Results
We begin by considering the filtering and equation learning errors for one-shot learning of Burgers’. These results are show in the top row of Figure 3(A) where we are reporting the errors for a fixed smooth initial condition with different number of observation points. Note, these errors were computed over the training data but they are different from the training error of the algorithm as they compare the filtered solution and the learned PDE on the test mesh $Y_{\text{test}}$. Since the observation points are random we are reporting the average errors along with the best and worst errors over different runs giving the shaded regions. We found that 1-step KEqL leads to the best filtering errors with a wider performance gap when $N_{\mathcal{Y}}$ is small. Interestingly, the performance gap appears to be more pronounced in the equation learning case where we observed an order of magnitude improvement over SINDy and almost two orders improvement over PINN-SR. We observed significantly larger errors for PINN-SR here which we attribute to difficulties in tuning and optimizing PINN models. In fact, we could often bring these errors down by hand tuning the algorithm in each instance of the experiments but automatic strategies proposed in [21] appeared to be very sensitive to the location of observation points.
This issue was exacerbated further when experiments were repeated with random initial conditions as depicted in Figure 3(B). Since both the observation points and initial conditions are chosen randomly, we also report the average errors over the experiments along with the best and worst errors across multiple runs, shown as shaded areas. A similar conclusion can be drawn here as in Figure 3(A), where the 1-step KEqL method continues to achieve the lowest filtering and equation learning errors, with a larger performance gap when $N_{\mathcal{Y}}$ is small. However, we now observe that the KEqL and SINDy methods exhibit greater robustness across different experiments, whereas the performance of the PINN-SR method deteriorates significantly. This decline is reflected in the increased variance in the results and the larger average error. We attribute this low performance to the use of the same model hyperparameters across different experiments. The main takeaway here is that KEqL and SINDy appear to be very robust to the choice of hyperparameters where as PINN-SR needs tailored hyperparameter tuning for every single run.
In Figure 3(C, D) we show two enlightening examples that show the difference in performance of 1-step methods vs 2-step. In (C) we hand picked an initial condition that leads to a solution with multiple (near) shocks. We see that despite scarcity of observations, 1-step KEqL and PINN-SR are able to capture the general shape of the solution while 2-step methods are significantly worse. In (D) we show an instance where we only observed the value of the solution on the boundary of the domain while in row (E) we solved the learned equation for a new initial condition. Once again we see that 1-step methods give better performance for filtering, however, when solving the learned PDE a noticeable performance gap appears across the three methods with 1-step KEqL giving the best solution. This example further reflects the performance gaps observed in panel (A), with the equation learning performance gaps being more pronounced compared to filtering.
@@
[…]
Figure 3. Representative numerical results for Burgers’ PDE (19): (A) The filtering and equation learning errors computed for the training functions for 1-step KEqL, SINDy, and PINN-SR using $M=1$ training pairs with different number of interior observations $N_{\mathcal{Y}}$; (B) Similar experiment as panel (A) but with randomized initial conditions; (C) An example application for an initial condition that leads to multiple shocks with scarce observations depicting the quality of filtering obtained using 1-step and 2-step methods; (D) Similar setup to row (C) with a smooth solution that is only observed on the boundary; (E) Depicting the solution to the PDEs that were learned in row (D) for a new initial condition.
@@ 6. Discussion
In conclusion, we presented the 1- and 2-step formulations of KEqL as an algorithm for learning nonlinear PDEs as well as their solution operators and filtering of observed states in scarce data regimes. Our theoretical results provided quantitative error bounds and convergence rates for our algorithms while our numerical experiments demonstrated significant gains in accuracy and robustness in comparison to existing methods in the literature. Most notably, our methods appear to be a lot less sensitive to choice of hyperparameters and demonstrate more consistent performance compared to equivalent neural net methods. Additionally, our exposition unifies various problems of interest to scientific machine learning under the same umbrella, i.e., equation learning, operator learning, and PDE solvers, all viewed as optimal recovery problems.
@@
JFOR, AH and BH were supported by the National Science Foundations grants DMS-2208535 ”Machine Learning for Inverse Problems” and DMS-2337678 ”Gaussian Processes for Scientific Machine Learning: Theoretical Analysis and Computational Algorithms”. YJ and HO acknowledge support from the Air Force Office of Scientific Research under MURI awards number FA9550-20-1-0358 (Machine Learning and Physics-Based Modeling and Simulation), FOA-AFRL-AFOSR-2023-0004 (Mathematics of Digital Twins), the Department of Energy under award number DE-SC0023163 (SEA-CROGS: Scalable, Efficient, and Accelerated Causal Reasoning Operators, Graphs and Spikes for Earth and Embedded Systems), the National Science Foundations under award number 2425909 (Discovering the Law of Stress Transfer and Earthquake Dynamics in a Fault Network using a Computational Graph Discovery Approach), and the Jet Propulsion Laboratory PDRDF 24AW0133 (Jupiter’s Radiation Environment: Assimilating Data with ML-driven Approaches). Additionally, HO acknowledges support from the DoD Vannevar Bush Faculty Fellowship Program.
@@ B.4.2. Representer theorem for 1-step KEqL (Proof of Theorem 2)
To simplify notation we first prove the theorem with $M=1$, i.e., with training data $(u(Y^{1}),f)$, and $\overline{P}=0$. At the end of the proof we comment on how the argument can be extended to the general case. Let us recall the corresponding optimal recovery problem for convenience:
@@
where $S=\Phi(v,Y).$
Recall that $Y$ is the dense set of collocation points that we use to impose the PDE constraint.
@@
$\widehat{u}(y)=\mathsf{U}(\phi,y)^{T}\widehat{\alpha},\quad\widehat{P}(s)=\mathsf{P}(S(\widehat{\alpha}),s)^{T}\widehat{\beta},$
@@
(32) $(\widehat{\alpha},\widehat{\beta})=$ $\operatorname*{arg\,min}_{\alpha\in\mathbb{R}^{QK},\>\beta\in\mathbb{R}^{K}}$ $\beta^{T}\mathsf{P}(S(\alpha),S(\alpha))\beta+\lambda_{1}\alpha^{T}\mathsf{U}(\phi,\phi)\alpha,$
s.t. $\mathsf{U}(\phi,Y^{1})^{T}\alpha=u(Y^{1}),\quad\text{and}\quad\mathsf{P}(S(\alpha),S(\alpha))^{T}\beta=f(Y),$
where $S(\alpha)=\Phi(\mathsf{U}(\phi,\cdot)^{T}\alpha,Y).$
Recall our notation $\mathsf{U}(\phi,Y^{m})\in\mathbb{R}^{QK\times N}$ for the matrix with columns $\mathsf{U}(\phi,y_{n})$, and $\mathsf{P}(S(\alpha),S(\alpha))\in\mathbb{R}^{K\times K}$ for the matrix with columns $\mathsf{P}(S(\alpha),s_{k})$ with $s_{k}=\Phi(\mathsf{U}(\phi,\cdot)^{T}\alpha,y_{k})$.
@@
For the general proof for an arbitrary number of functions $M$ the same idea follows since the equivalent form (33) now contains the slack tensor $Z\in\mathbb{R}^{K\times Q\times M}$ instead of the matrix $Z$ above since each $v^{m}$ will be represented by a matrix $Z^{m}$ in our proof. Additionally $S$ will then contain not only $K$ points but $MK$ points corresponding to each of the $v^{m}$ and their requisite partial derivatives evaluated at the collocation points. To obtain the result when $\overline{P}\neq 0$ the proof simply follows by replacing the data $f(Y)$ in (31) with $\overline{P}(S)$ and similarly in subsequent arguments in the proof.
@@ D.2.1. Experiments with increasing number of observations
Convergence history of the optimization algorithm for some sample cases of learning the equation is shown in Figure 10(A-B).
@@ D.2.3. A one-shot learning experiment
Figure 9. Quantitative values of relative filtering and operator learning errors for Burgers’ PDE Equation 19 with different ICs corresponding to the one-shot learning examples Ex. 1 (Section D.2.2) and Ex. 2 (Section D.2.3) in Figure 3(C-E).